# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

### v1

@@ Abstract
Graph clustering is a longstanding topic in machine learning. In recent years, deep learning methods have achieved encouraging results, but they still require predefined cluster numbers $K$, and typically struggle with imbalanced graphs, especially in identifying minority clusters. The limitations motivate us to study a challenging yet practical problem: deep graph clustering without $K$ considering the imbalance in reality. We approach this problem from a fresh perspective of information theory (i.e., structural information). In the literature, structural information has rarely been touched in deep clustering, and the classic definition falls short in its discrete formulation, neglecting node attributes and exhibiting prohibitive complexity. In this paper, we first establish a new Differentiable Structural Information, generalizing the discrete formalism to continuous realm, so that the optimal partitioning tree, revealing the cluster structure, can be created by the gradient backpropagation. Theoretically, we demonstrate its capability in clustering without requiring $K$ and identifying the minority clusters in imbalanced graphs, while reducing the time complexity to $O(N)$ w.r.t. the number of nodes. Subsequently, we present a novel IsoSEL framework for deep graph clustering, where we design a hyperbolic neural network to learn the partitioning tree in the Lorentz model of hyperbolic space, and further conduct Lorentz Tree Contrastive Learning with isometric augmentation. As a result, the partitioning tree incorporates node attributes via mutual information maximization, while the cluster assignment is refined by the proposed tree contrastive learning. Extensive experiments on five benchmark datasets show the IsoSEL outperforms $14$ recent baselines by an average of $+1.3\%$ in NMI.
Index Terms:
Graph Clustering, Deep Learning, Structural Entropy, Hyperbolic Space, Lorentz Group.
@@ I Introduction
  • The requirement of predefined cluster numbers $K$. So far, deep graph clustering cannot work without a predefined cluster number $K$, and thus one is required to correctly provide $K$ before clustering, which is often impractical in real cases. On the one hand, it is challenging to estimate the number of node clusters. The methods, e.g., Bayesian information criterion [7] and recent RGC [8], repeatedly train and test the deep model to search the optimal $K$, and the searching process typically results in high computation overhead. On the other hand, existing solutions without graph structures (e.g., Bayesian non-parametric methods [9], DBSCAN [10], and hierarchical clustering [11]) cannot be directly applied owing to the inter-correlation among nodes. That is, deep graph clustering without $K$ is largely underexplored.
@@
  • Clustering the imbalanced graphs. The imbalance is prevalent in real-world graphs, where a large proportion of vertices belong to the majority clusters, while minority clusters contain significantly fewer vertices [12, 13]. Take the popular Cora dataset [14] as an example: $26.8\%$ of the papers belong to Neural Network cluster, whereas only $4.8\%$ fall under Rule Learning cluster. Deep graph clustering models typically rely on the representation learning, and minority clusters tend to be underrepresented due to their limited sample sizes compared to the majority clusters [15]. This imbalance increases the risk of generating biased clustering boundaries that favor majority clusters, as evidenced by empirical observations in our Case Study. Also, the challenge becomes even greater when the number of clusters is unknown.
Figure 1: In IsoSEL, we learn the partitioning tree of structural information with the LSEnet in hyperbolic space, and then propose Lorentz tree contrastive learning to refine the partitioning tree for graph clustering without $K$.
In this paper, we revisit deep graph clustering with an unknown $K$, considering the imbalance in real-world graphs, and approach this problem from a fresh perspective based on information theory. Stemming from Shannon entropy, structural information [16] has been proposed to measure the uncertainty in graphs. According to structural information minimization, the optimal partitioning tree can be constructed to describe a graph’s self-organization without requiring $K$. While this sheds light on the targeted problem, it presents several significant challenges to deep clustering: (1) Connection to Clustering. Structural information has recently garnered increasing research attention [17, 18, 19], but its theoretical connection to clustering remains unclear. As a result, structural information has not yet been introduced to graph clustering. (2) Discrete Formulation. The discrete nature of structural information prevents gradient backpropagation, posing a major challenge in training deep graph models. (3) Attribute Ignorance. Structural information operates solely on the graph structure overlooking node attributes, which are often equally important to graph clustering. (4) Computational Complexity. Constructing a partitioning tree of height $3$ yields the complexity of $O(\lvert N\rvert^{3}\log^{2}\lvert N\rvert)$, where $N$ is the number of nodes. Note that, the complexity scales exponentially with tree height [16], making it infeasible for real-world graphs.
@@
To tackle these challenges, we establish a new differentiable structural information (DSI), generalizing the discrete formalism to the continuous realm. In DSI, we reformulate the structural information with level-wise assignment matrices. We prove that DSI upper bounds the graph conductance, a well-defined clustering objective, and thus DSI emerges as a new differentiable graph clustering objective without requiring the number of clusters $K$. Thus we create a deep partitioning tree $\mathcal{T}_{\operatorname{net}}$ parametrized by a neural network. Different from the partitioning tree of [16] constructed by heuristic algorithm, the proposed deep tree can be learned by gradient backpropagation according to DSI minimization. Theoretically, we show that the structural entropy of the optimal $\mathcal{T}_{\operatorname{net}}$ well approximates that of [16] under slight constraint. Next, we reduce the time complexity to the linear $\mathcal{O}(N)$ with respect to the number of nodes by exploring the equivalence in level-wise assignments. Furthermore, we discover another advantage of the deep partitioning tree that it is able to identify the minority clusters from the majority ones, which is suitable for clustering real graphs that often exhibit imbalanced cluster distributions.
Grounded on DSI theory, we present a novel Isometric Structural Entropy Learning (referred to as IsoSEL) to learn and refine the deep partitioning tree in the hyperbolic space for graph clustering without $K$. The overall architecture is illustrated in Fig. 1, consisting of the Lorentz Structural Entropy net (LSEnet) from the conference version [^1] The preliminary results have been published in the proceedings of the 41st International Conference on Machine Learning (ICML) as an Oral Paper [20]., and Lorentz tree contrastive learning proposed in this full version. As in the part B of Fig. 1, LSEnet is designed to parameterize the deep partitioning tree in the Lorentz model of hyperbolic space. The novelty lies in exploring the alignment between hyperbolic space and hierarchical/tree-like structures. Specifically, we first embed the leaf nodes of the partitioning tree using Lorentz convolution over the graph. Then, we recursively learn the parent nodes from the bottom (leaf nodes) to the top (root node) using the proposed Lorentz assigner.
In IsoSEL, another key innovation is the proposed Lorentz Tree Contrastive Learning (as shown in part C of Fig. 1), which refines the deep partitioning tree in the hyperbolic space. To achieve this, we first formulate the isometric augmentation to generate an augmented graph, integrating attribute similarity into structural information from the representation perspective using the notion of Gauss curvature. Second, we design a projection head for tree contrastive learning. Unlike graph contrastive learning, the main challenge in contrasting trees lies in ensuring consistency between leaf and parent nodes in hyperbolic space, given the unified nature of the partitioning tree. To address this, we instantiate a neural Lorentz group in hyperbolic space. Specifically, we parameterize the transformation in the Lorentz group using a closed form matrix-vector multiplication, which supports gradient descent optimization. This parameterization ensures that the projection head preserves the hyperbolic manifold and maintains consistency between leaf and parent nodes, leveraging the linearity of the Lorentz group. Finally, we utilize the Minkowski inner product to contrast the deep partitioning trees, refining the hyperbolic representations and enhancing cluster assignment. As a result, IsoSEL seamlessly integrates the strengths of both structural entropy and hyperbolic space for graph clustering.
@@
  • Theory. We establish the differentiable structural information (DSI) in continuous realm, so that a deep partitioning tree with linear computational complexity is created by gradient backpropagation for graph clustering without $K$.
  • Neural Architecture. We present a novel IsoSEL approach where the hyperbolic LSEnet is designed to learn the deep partitioning tree with Lorentz convolution and assigner, connecting DSI to hyperbolic geometry.
  • Learning Paradigm. We propose a Lorentz tree contrastive learning with the isometric augmentation to refine the deep partitioning tree in the hyperbolic space, while incorporating the attribute information.
  • Experiment. Extensive real world results show the superiority of IsoSEL for graph clustering, and we examine that biased boundary among imbalanced clusters is alleviated with case study and visualization.
@@ II-2 Structural Information & Structural Entropy
$\mathcal{H}^{1}(G)=-\sum_{v\in\mathcal{V}}\frac{d_{v}}{\operatorname{Vol}(G)}\log_{2}\frac{d_{v}}{\operatorname{Vol}(G)},\vspace{-0.03in}$ (1)
yielding as a constant for certain given graph. The $H$-dimensional structural information is accompanied with a partitioning tree $\mathcal{T}$ of height $H$. Concretely, in $\mathcal{T}$ with root node $\lambda$, each tree node $\alpha$ is associated with a subset of $\mathcal{V}$, referred to as module $T_{\alpha}\subset\mathcal{V}$ [^2] The vertex of partitioning tree is termed as “tree node” to distinguish from the node in graphs. and its immediate predecessor is written as $\alpha^{-}$. The leaf node of the tree corresponds to the node of graph $G$. Accordingly, the module of the leaf node is a singleton, while $T_{\lambda}$ is the node set $\mathcal{V}$. The structural information assigned to each non-root node $\alpha$ is defined as
$\mathcal{H}^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\operatorname{Vol}(G)}\log_{2}\frac{V_{\alpha}}{V_{\alpha^{-}}},\vspace{-0.03in}$ (2)
@@
$\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{\alpha\in\mathcal{T},\alpha\neq\lambda}\mathcal{H}^{\mathcal{T}}(G;\alpha).\vspace{-0.03in}$ (3)
@@
$\mathcal{H}^{H}(G)=\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(G),\quad\mathcal{T}^{*}=\arg_{\mathcal{T}}\min\mathcal{H}^{\mathcal{T}}(G),\vspace{-0.03in}$ (4)
@@ II-3 Hyperbolic Space & Lorentz Group
$\mathbb{L}^{\kappa,d}=\{\boldsymbol{x}\in\mathbb{R}^{d+1}:\langle\boldsymbol{x},\boldsymbol{x}\rangle_{\mathbb{L}}=\frac{1}{\kappa},x_{0}>0\},\vspace{-0.03in}$ (5)
@@
$\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}=\boldsymbol{x}\mathbf{R}\boldsymbol{y},\quad\mathbf{R}=\operatorname{diag}(-1,1,\cdots,1),\vspace{-0.03in}$ (6)
@@
In group theory, the Lorentz group studies the Lorentz model of hyperbolic space. To be specific, Lorentz group is a (generalized) orthogonal transformation group that collects the Lorentz transformation, preserving the isometry regarding the Minkowski metric. With polar decomposition, any Lorentz transformation can be decomposed into a combination of Lorentz boost and Lorentz rotation [21, 22]. Both of them are given as special orthogonal matrices, posing a significant challenge to parameterization. Please refer to Appendix B for further facts.
@@ II-4 Notations
In this paper, we introduce the novel Differentiable Structural Information (DSI) and present Isometric Structural Entropy Learning (IsoSEL) for deep graph clustering. The key innovation of DSI lies in constructing a deep partitioning tree in hyperbolic space through gradient backpropagation. Additionally, we propose Lorentz tree contrastive learning with theoretical guarantees to further refine the deep partitioning tree, enhancing cluster assignment.
@@ III Differentiable Structural Information
III-A New Formulation
The structural information in Eq. (1) cannot be optimized via gradient-based methods. To bridge this gap, we present a new formulation of structural information with learnable level-wise assignment, which is shown to be equivalent to the classic formulation in Eq. (3).
@@ III-A New Formulation
[…]
@@
$\mathcal{H}^{\mathcal{T}}(G;h)=-\frac{1}{V}\sum\limits_{k=1}^{N_{h}}(V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij})\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}\vspace{-0.03in}$ (7)
@@
$V^{h}_{k}$ $=\sum\nolimits_{i=1}^{N}S^{h}_{ik}d_{i},\quad V^{h-1}_{k^{-}}=\sum\nolimits_{k^{\prime}=1}^{N_{h-1}}\mathbf{C}_{kk^{\prime}}^{h}V^{h-1}_{k^{\prime}}.\vspace{-0.03in}$ (9)
@@ III-B Connecting Structural Information and Graph Clustering
$\mathcal{H}^{1}(G)=\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}}),\vspace{-0.03in}$ (10)
@@
Theorem III.5 (Connection Graph Clustering).

### v2

# ASIL: Augmented Structural Information Learning for Deep Graph Clustering in Hyperbolic Space

@@ Abstract
Graph clustering is a longstanding topic in machine learning. In recent years, deep learning methods have achieved encouraging results, but they still require predefined cluster numbers $K$, and typically struggle with imbalanced graphs, especially in identifying minority clusters. The limitations motivate us to study a challenging yet practical problem: deep graph clustering without $K$ considering the imbalance in reality. We approach this problem from a fresh perspective of information theory (i.e., structural information). In the literature, structural information has rarely been touched in deep clustering, and the classic definition falls short in its discrete formulation, neglecting node attributes and exhibiting prohibitive complexity. In this paper, we first establish a differentiable structural information, generalizing the discrete formalism to continuous realm, so that we design a hyperbolic deep model (LSEnet) to learn the neural partitioning tree in the Lorentz model of hyperbolic space. Theoretically, we demonstrate its capability in clustering without requiring $K$ and identifying minority clusters in imbalanced graphs. Second, we refine hyperbolic representations of the partitioning tree, enhancing graph semantics, for better clustering. Contrastive learning for tree structures is non-trivial and costs quadratic complexity. Instead, we further advance our theory by discovering an interesting fact that structural entropy indeed bounds the tree contrastive loss. Finally, with an efficient reformulation, we approach graph clustering through a novel augmented structural information learning (ASIL), which offers a simple yet effective objective of augmented structural entropy to seamlessly integrates hyperbolic partitioning tree construction and contrastive learning. With a provable improvement in graph conductance, ASIL achieves effective debiased graph clustering in linear complexity with respect to the graph size. Extensive experiments show the ASIL outperforms $20$ strong baselines by an average of $12.42\%$ in NMI on Citeseer dataset.
@@ I Introduction
Challenge 1: The requirement of predefined cluster numbers $K$. So far, deep graph clustering typically relies on predefined cluster numbers K, and thus one is required to provide an accurate K prior to clustering, which is often impractical in real cases. On the one hand, it is challenging to estimate the number of node clusters. The methods, e.g., Bayesian information criterion [49] and recent RGC [34], repeatedly train and test the deep model to search for the optimal $K$, and the searching process typically results in high computation overhead. On the other hand, existing solutions without graph structures (e.g., Bayesian non-parametric methods [18], DBSCAN [16], and hierarchical clustering [10]) cannot be directly applied owing to the inter-correlation among nodes. Thus, deep graph clustering without $K$ remains largely underexplored.
@@
Challenge 2: Clustering the imbalanced graphs. The imbalance is prevalent in real-world graphs, where a large proportion of vertices belong to the majority clusters, while minority clusters contain significantly fewer vertices [23, 39]. Take the popular Cora dataset [50] as an example: $26.8\%$ of the papers belong to Neural Network cluster, whereas only $4.8\%$ fall under Rule Learning cluster. Deep graph clustering models typically rely on the representation learning, and minority clusters tend to be underrepresented due to their limited sample sizes compared to the majority clusters [53]. This imbalance increases the risk of generating biased clustering boundaries that favor majority clusters, as evidenced by empirical observations in our Case Study. Also, the challenge becomes even greater when the number of clusters is unknown.
Figure 1: Illustration of ASIL. With a novel objective of augmented structural entropy, ASIL seamlessly integrates partitioning tree construction and contrastive learning in the Lorentz model of hyperbolic space. It achieves debiased node clustering in linear complexity without requiring the number of clusters.
[…]
@@
To tackle these challenges, we establish a new differentiable structural information (DSI), generalizing the discrete formalism to the continuous realm. In DSI, we reformulate the structural information with level-wise assignment matrices. We prove that DSI upper bounds the graph conductance, a well-defined clustering objective, and thus DSI emerges as a new differentiable graph clustering objective without requiring the number of clusters $K$. Thus we create a deep partitioning tree $\mathcal{T}_{\operatorname{net}}$ parametrized by a neural network. Different from the partitioning tree of [27] constructed by heuristic algorithm, the proposed deep tree can be learned by gradient backpropagation according to DSI minimization. Theoretically, we show that the structural entropy of the optimal $\mathcal{T}_{\operatorname{net}}$ well approximates that of [27] under slight constraint. it is able to identify the minority clusters from the majority ones, which is suitable for clustering real graphs that often exhibit imbalanced cluster distributions. Furthermore, we uncover another advantage of the deep partitioning tree—its capability to distinguish minority clusters from majority ones, making it suitable for clustering real-world graphs that often exhibit imbalanced distributions.
[…]
Going beyond, in this full version, we aim to enhance both effectiveness and efficiency of structural entropy clustering. For effectiveness, we refine the partitioning tree representation in the Lorentz model of hyperbolic space via contrastive learning. Different from graph contrastive learning, we derive a neural Lorentz boost (Theorem VI.1, VI.2) as a special projection head of consistency insurance—pushing leaf nodes together with the parent hierarchies as a unity—and formulate a tree contrastive loss considering partitioning tree branches for clustering assignment. However, its quadratic complexity prevents the clustering model scaling up. For efficiency, we further advance our established theory by an interesting new finding (Theorem VII.1) that structural entropy indeed bounds the contrastive loss. Accordingly, we connect the entropy (of structural perspective) and contrastive learning (of representational perspective) through a novel virtual graph constructed from the representations of the proposed Lorentz boost, maintaining consistency insurance, so that optimizing the virtual graph implicitly performs the tree contrastive learning, as shown in Fig. 1B. Meanwhile, we equivalently reformulate structural entropy by exploring level-wise assignments, achieving linear time complexity. Finally, we propose an expressive objective of augmented structural entropy, fusing the virtual graph described above, which seamlessly integrates partitioning tree construction and contrastive learning. With a provable improvement in graph conductance (Theorem VII.2), ASIL achieves effective debiased graph clustering in linear complexity with respect to the graph size.
@@
  • Theory. We establish the differentiable structural information in continuous realm, so that a deep partitioning tree with linear computational complexity is created by gradient backpropagation for graph clustering without $K$. We further connect the theory of structural information to constrastive learning by proving a new bound.
  • Neural Architecture. We present a hyperbolic clustering architecture that first fuses a parameterized virtual graph and then leverages LSEnet to learn the partitioning tree in the Lorentz model of hyperbolic space.
  • Learning Paradigm. We propose a novel Augmented Structural Information Learning that integrates partitioning tree construction and contrastive learning for more effectively and efficiently debiased graph clustering without requiring $K$.
  • Experiment. Extensive real world results show the superiority of ASIL for graph clustering, and we examine that biased boundary among imbalanced clusters is alleviated with case study and visualization.
@@ II-2 Structural Information & Structural Entropy
$\mathcal{H}^{1}(G)=-\sum_{v\in\mathcal{V}}\frac{d_{v}}{\operatorname{Vol}(G)}\log_{2}\frac{d_{v}}{\operatorname{Vol}(G)},\vskip-2.168pt$ (1)
yielding as a constant for certain given graph. The $H$-dimensional structural information is accompanied with a partitioning tree $\mathcal{T}$ of height $H$. Concretely, in $\mathcal{T}$ with root node $\lambda$, each tree node $\alpha$ is associated with a subset of $\mathcal{V}$, referred to as module $T_{\alpha}\subset\mathcal{V}$. [^3] The vertex of partitioning tree is termed as “tree node” to distinguish from the node in graphs. and its immediate predecessor is written as $\alpha^{-}$. The leaf node of the tree corresponds to the node of graph $G$. Accordingly, the module of the leaf node is a singleton, while $T_{\lambda}$ is the node set $\mathcal{V}$. The structural information assigned to each non-root node $\alpha$ is defined as
$\mathcal{H}^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\operatorname{Vol}(G)}\log_{2}\frac{V_{\alpha}}{V_{\alpha^{-}}},\vskip-2.168pt$ (2)
@@
$\mathcal{H}^{\mathcal{T}}(G)=\sum\nolimits_{\alpha\in\mathcal{T},\alpha\neq\lambda}\mathcal{H}^{\mathcal{T}}(G;\alpha).\vskip-2.168pt$ (3)
@@
$\mathcal{H}^{H}(G)=\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(G),\quad\mathcal{T}^{*}=\arg_{\mathcal{T}}\min\mathcal{H}^{\mathcal{T}}(G),\vskip-2.168pt$ (4)
@@ II-3 Hyperbolic Space & Lorentz Group
$\mathbb{L}^{\kappa,d}=\{\boldsymbol{x}\in\mathbb{R}^{d+1}:\langle\boldsymbol{x},\boldsymbol{x}\rangle_{\mathbb{L}}=\frac{1}{\kappa},x_{0}>0\},\vskip-5.05942pt$ (5)
@@
$\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}=\boldsymbol{x}\mathbf{R}\boldsymbol{y},\quad\mathbf{R}=\operatorname{diag}(-1,1,\cdots,1),\vskip-5.05942pt$ (6)
@@
In group theory, the Lorentz group studies the Lorentz model of hyperbolic space. To be specific, Lorentz group is a (generalized) orthogonal transformation group that collects the Lorentz transformation, consisting of Lorentz boosts and Lorentz rotations, which preserves the isometry regarding the Minkowski metric [9, 47]. Both of them are given as special orthogonal matrices, posing a significant challenge to parameterization. Further facts on hyperbolic space and Lorentz group are provided in Appendix V.
@@ II-4 Notations
In this paper, we establish the theory of differential structural information, demonstrating its capacity in graph clustering and connecting it to contrastive learning. Thus, we approach graph clustering through Augmented Structural Entropy Learning (ASIL), whose key innovation lies in the integration of partitioning tree construction and contrastive learning. Without requiring the number of clusters, ASIL effectively achieves debiased node clustering in linear complexity with respect to the graph size.
@@ III Differentiable Structural Information
III-A New Formulation: A Differentiable Equivalence
We introduce a differentiable version of structural entropy where the hard indicator for calculating the volume is relaxed as soft level-wise assignments. It is shown to be equivalent to the classic formulation, while enabling the optimization via gradient backpropagation.
@@ III-A New Formulation: A Differentiable Equivalence
For a partitioning tree $\mathcal{T}$ with height $H$, assuming that the number of tree nodes at the $h$-th level is $N_{h}$, we define a level-wise parent assignment matrix $\mathbf{C}^{h}\in\{0,1\}^{N_{h}\times N_{h-1}}$ from $h$-th to (h1)(h-1)-th level, where $\mathbf{C}^{h}_{ij}=1$ means the $i$-th node of $\mathcal{T}$ at $h$-th level is the child node of $j$-th node at (h1)(h-1)-th level.
@@
$\mathcal{H}^{\mathcal{T}}(G;h)=-\frac{1}{V}\sum\limits_{k=1}^{N_{h}}(V^{h}_{k}-\sum\limits_{(i,j)\in\mathcal{E}}S^{h}_{ik}S^{h}_{jk}w_{ij})\log_{2}\frac{V^{h}_{k}}{V^{h-1}_{k^{-}}}\vskip-2.168pt$ (7)
@@
$V^{h}_{k}$ $=\sum\nolimits_{i=1}^{N}S^{h}_{ik}d_{i},\quad V^{h-1}_{k^{-}}=\sum\nolimits_{k^{\prime}=1}^{N_{h-1}}\mathbf{C}_{kk^{\prime}}^{h}V^{h-1}_{k^{\prime}}.\vskip-2.168pt$ (9)
@@ III-B Connecting Structural Information and Graph Clustering
$\mathcal{H}^{1}(G)=\sum_{h=1}^{H}\sum_{j=1}^{N_{h-1}}\frac{V^{h-1}_{j}}{V}E([\frac{C^{h}_{kj}V^{h}_{k}}{V^{h-1}_{j}}]_{k=1,...,N_{h}}),\vskip-2.168pt$ (10)
@@
Theorem III.5 (Connection to Graph Clustering).
@@
Remark. Different from classical graph entropies measuring the uncertainty on graph, structural entropy also models the hierarchical self-organization of graphs through a partitioning tree. This enables it to serve as a clustering objective without requiring the number of clusters, a property absent in traditional graph entropy measures.